# The LHAASO Collaboration

### v2

@@ Abstract
The diffuse Galactic gamma-ray emission is a very important tool used to study the propagation and interaction of cosmic rays in the Milky Way. In this work, we report the measurements of the diffuse emission from the Galactic plane, covering Galactic longitudes from $15^{\circ}$ to $235^{\circ}$ and latitudes from $-5^{\circ}$ to $+5^{\circ}$, in an energy range of 1 TeV to 25 TeV, with the Water Cherenkov Detector Array (WCDA) of the Large High Altitude Air Shower Observatory (LHAASO). After masking the sky regions of known sources, the diffuse emission is detected with $24.6\sigma$ and $9.1\sigma$ significance in the inner Galactic plane($15^{\circ}<l<125^{\circ}$, $|b|<5^{\circ}$) and outer Galactic plane ($125^{\circ}<l<235^{\circ}$, $|b|<5^{\circ}$), respectively. The WCDA spectra in both regions can be well described by a power-law function, with spectral indices of $-2.67\pm 0.05_{\rm stat}$ in the inner region and $-2.83\pm 0.19_{\rm stat}$ in the outer region, respectively. Combined with the Square Kilometer Array (KM2A) measurements at higher energies, a clear softening of the spectrum is found in the inner region, with change of spectral indices by $\sim 0.5$ at a break energy around $30$ TeV. The fluxes of the diffuse emission are higher by a factor of $1.5-2.7$ than the model prediction assuming local CR spectra and the gas column density, which are consistent with those measured by the KM2A. Along Galactic longitude, the spatial distribution of the diffuse emission shows deviation from that of the gas column density. The spectral shape of the diffuse emission are possibly variation in different longitude region. The WCDA measurements bridge the gap between the low-energy measurements by space detectors and the ultra-high-energy observations by LHAASO-KM2A and other experiments. The results indicate further that the conventional model prediction of diffuse emission from cosmic ray interactions is not enough to account for the observations.
@@
Introduction. — One of the most important unresolved problems in astrophysics is the origin and propagation of cosmic rays (CRs). Besides the direct measurements of energy spectra and arrival directions of individual composition of CRs, the diffuse $\gamma$-ray emission plays a unique and complementary role in constraining the origin and propagation of CRs. The diffuse $\gamma$-ray emission is usually expected to be produced through interactions between CRs (nuclei and electrons/positrons) and the interstellar medium (ISM) or radiation field [1, 2, 3]. Modeling of the diffuse emission in the high-energy $\gamma$-ray bands (HE, $\lesssim$0.1 TeV) is in general consistent with the all-sky measurements by Fermi Large Area Telescope [4], supporting the basic framework of CR propagation and interaction in the Milky Way.
@@
We also derive the spatial distributions of the WCDA diffuse emission along Galactic latitudes and longitudes, as shown in Fig. 4. The results are compared with the gas distributions obtained from the sum of atomic hydrogen from the HI4PI survey [29], molecular hydrogen from the CO survey [30] with a CO-to-H2 conversion factor of $X_{\rm CO}=1.6\times 10^{20}$ cm2 $({\rm K\,km\,s^{-1}})^{-1}$ [4], and an ionized hydrogen component from the model of [31]. It is shown that the latitude distributions are roughly consistent with the gas distributions in both the inner and outer regions. The fitting of the gas template to the data gives $p$-values of $0.003-0.013$ of the two regions. However, the longitude distribution deviates clearly from the gas distribution. The gas template fitting shows a $p$-value of $6.0\times 10^{-21}$. Similar results have also been obtained for the KM2A measurements at higher energies [11] (see Fig. S6 of the Supplemental Material for the comparison between WCDA and KM2A).
@@
Systematic uncertainties. — One of the major sources of systematic uncertainties is the background estimate. We vary the time window from ±$\pm 5$ hours to ±$\pm 12$ hours, and also change the mask regions for background estimate to $|b|\leq 10^{\circ}$ for all declinations to obtain the variations of the results. The other effect related with the background is the spurious large-scale efficiency correction (see Sec. B of the Supplemental Material), for which we adopt different smoothing angles of $15^{\circ}$, $20^{\circ}$, $25^{\circ}$, and $30^{\circ}$. We find that the impacts of those changes on parameters $\phi_{0}$ are about $6\%~{}(16\%)$ and on $\alpha$ are about 0.05 (0.13), for the inner (outer) region. The variation of the atmosphere condition during the operation affects the detection efficiency, which is not properly modelled in the simulation. For point-like sources this effect is estimated to be about 8% for the flux ($\phi_{0}$) and 0.02 for the spectral index ($\alpha$) [19]. Similar impacts on the diffuse emission is expected. Combining these results, we obtain the total systematic uncertainties as 10% (18%) for $\phi_{0}$ and 0.05 (0.13) for $\alpha$, for the inner (outer) region, respectively. The discussion of systematic uncertainties of the KM2A results can be referred to Ref. [11]. The systematic uncertainty of the flux in each energy bin has been shown in Fig. 2 and also in Tables S2 and S3 of the Supplemental Material.
@@
Discussion. — We compare the data with the predictions of the hadronic interactions between CRs and the ISM. We adopt the parameterizations of the local proton and helium spectra as given in Ref. [11], which include the uncertainties of the current measurements, especially for those measured by indirect detection experiments above $\sim 100$ TeV. We assume that the CR intensity is uniform in the Galaxy, and calculate the $\gamma$-ray emission using the local CR spectra and the line-of-sight integrated gas content. The $\gamma$-ray production cross section used is from the AAfrag package [32]. The expected diffuse $\gamma$-ray fluxes are shown by grey shaded bands in Fig. 2. We can see that in the inner region the measurements are higher by a factor of $\sim 2.7$ than the prediction, and in the outer region the measurements are higher by $\sim 1.5$ times. These results are consistent with those found in the LHAASO-KM2A analysis [11]. The HAWC measurements in a different sky region of the inner Galaxy also showed a factor of 2 higher fluxes compared with a DRAGON model prediction [10]. The spectral shape of the prediction is slightly different from the measurement in the inner region from tens of TeV to PeV. However, in the outer region, the predicted spectrum is consistent with the data within uncertainties. We point out that the observed spectral break around 30 TeV in $\gamma$ rays may not correspond to the knee of CRs which is expected to break in hundreds of TeV given that the current measured spectra of protons and helium show the knee feature around several PeV [33, 34]. More data are needed to clarify whether there are imprints of the knee feature in the diffuse $\gamma$-ray spectra.
@@
The break may be due to the addition of an unresolved source population whose spectrum show a cutoff around tens of TeV. If we compare the data (including those measured by Fermi-LAT between 1 and 500 GeV [12], updated with the new mask of this work) with the model prediction111 The prediction is based on a GALPROP model, which assumes a cylindrical propagation halo of the Milky Way, employs a numerical method to solve the transport equations, and includes experimental/observational inputs about the nuclear cross sections and gas distributions [2]. The propagation and source injection parameters were adjusted to fit the up-to-date measurements of CR primary and secondary nuclei [35]. Note that, a homogeneous, isotropic diffusion is assumed when calculating the CR propagation, which may be over-simplified [36]. in a wider energy range, from GeV to PeV, we find that the excess mainly exists from several GeV to tens of TeV (see Fig. 5). Such an excess may come from a population of unresolved, low surface brightness sources such pulsar wind nebulae or pulsar halos [1, 37, 38, 13, 16]. The spatial variations of the spectra of the diffuse emission may further supports the hypothesis that unresolved sources may exist, although other possibilities such as spatially-dependent spectral slopes of CRs can not be excluded. Following Ref. [12], we add an empirical component with spectrum $\propto E^{-2.45}\exp\left(-E/20\,{\rm TeV}\right)$ to the CR propagation model prediction, and find a good agreement with the wide-band measurements, as shown in Fig. 5. Other sources like young massive star clusters [39], the confinement and interaction of CRs around acceleration sources [40, 41, 42], as well as the change of the canonical propagation scenario [43, 44, 45, 46] were also discussed in literature. The multi-messenger analysis of the diffuse $\gamma$ rays and neutrinos [28] would be helpful in understanding the nature of the diffuse emission.
@@
Summary. — In this work we measure the diffuse $\gamma$-ray emission in an energy range of about $1-25$ TeV from the Galactic plane using the LHAASO-WCDA data. The detection significance of the diffuse emission, after masking sky regions around resolved sources, is $24.6\sigma$ for the inner Galactic plane ($15^{\circ}<l<125^{\circ}$, $|b|<5^{\circ}$) and $9.1\sigma$ for the outer Galactic plane ($125^{\circ}<l<235^{\circ}$, $|b|<5^{\circ}$) regions, respectively. The diffuse emission from the outer Galactic plane is for the first time detected in the multi-TeV energy range. The WCDA measurements fill the gap between direct detection by space telescopes and ultra-high-energy measurements by groundbased experiments, and offer a full energy coverage from GeV to PeV for the Galactic diffuse $\gamma$-ray emission. The SEDs of the diffuse emission in the two regions have been measured with high precision. The WCDA spectra are consistent with power-law models, with indices of $-2.67\pm 0.05_{\rm stat}$ and $-2.83\pm 0.19_{\rm stat}$. The joint WCDA-KM2A spectrum shows a break around 30 TeV in the inner region, with indices changing by about $0.5$. The break in the outer region is not significant. We also show that there are slight spectral variations across the Galactic plane in the inner region. The measured fluxes are higher by a factor of $1.5-2.7$ than the simple prediction of hadronic interactions between CRs and the ISM, indicating that there are unmodelled source components or spatial variations of CR distribution. The former is preferred if the low-energy measurements by Fermi-LAT and high-energy measurements by LHAASO-KM2A are taken into account.
@@
Acknowledgements. — We would like to thank all staff members who work at the LHAASO site above 4400 meter above the sea level year-round to maintain the detector and keep the water recycling system, electricity power supply and other components of the experiment operating smoothly. We are grateful to Chengdu Management Committee of Tianfu New Area for the constant financial support for research with LHAASO data. We appreciate the computing and data service support provided by the National High Energy Physics Data Center for the data analysis in this paper. This work is supported by the following grants: the National Natural Science Foundation of China (Nos. 12220101003, 12173039, 12393851, 12393852, 12393853, 12393854, 12205314, 12105301, 12305120, 12261160362, 12105294, U1931201, 12375107), the Department of Science and Technology of Sichuan Province, China (No. 24NSFSC2319), the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061), the National Science and Technology Development Agency (NSTDA) of Thailand, and the National Research Council of Thailand (NRCT) under the High-Potential Research Team Grant Program (N42A650868).
Supplemental Material
.1 Correction of large-scale structures
The background estimate method may introduce some large-scale structures due to the efficiency variation in a relatively long time window (10 hours). Similar phenomena have also been observed in the KM2A analysis [11]. We adopt the same method as that used in the KM2A analysis to correct these structures, via smoothing the ON and BKG counts map (after masking) with top-hat kernels, and then correct the background by multiplying smoothing $N_{\rm on}/N_{\rm BKG}$ of each pixel. The benchmark radius of the smooth kernel is $20^{\circ}$, and other radii of $15^{\circ}$, $25^{\circ}$, $30^{\circ}$ have been adopted to estimate the systematic uncertainty due to this correction. The significance maps before and after the correction of large-scale structures are shown in Fig. S1.
Figure S1: Significance maps in celestial coordinate for $N_{\rm hit}\geq 100$, before (top) and after (bottom) the large-scale structure correction. The black solid lines show the ROIs of this analysis.
.2 Residual contamination from resolved sources
There could be residual emission from known sources whose morphologies are assumed to be Gaussian, given the current mask radius of $2.5\sigma$. To estimate the contamination to diffuse emission from known sources, we divide the sky region into different rings around each source, with radius of each ring being a factor of $n$ times of the source extension $\sigma$. The top panel of Fig. S2 illustrates the division of the sky. We construct two templates to fit the spatial distribution of events. One is the diffuse template from the gas distribution, and the other is the source template from all resolved sources weighted by their fluxes in corresponding energy band. The fitting result in one energy bin for the inner region is shown in the bottom panel of Fig. S2. The residual contamination can be obtained by setting $n=2.5$ as adopted in the analysis, which is given in Table S1. Such residual source contributions have been subtracted when deriving the diffuse emission fluxes.
Figure S2: Top: Schematic plot of sky regions of rings surrounding resolved sources. Bottom: Fitting results of the simulated signal intensities as functions of multiplicative factor $n$ of the source extension $\sigma$, for resolved sources and diffuse emission.
Figure S3: One-dimensional significance distributions for $N_{\rm hit}\geq 200$ (corresponding to energies bigger than several TeV) after masking sources. Top two panels are for our ROIs, and bottom panels are the reference sky regions shifted to higher Galactic latitudes by 20 degrees.
Table S1: Residual fractions of resolved sources to the total signal in different hit bins.
$N_{\rm hit}$ Inner Galaxy (%) Outer Galaxy (%)
60-100 $11.12\pm 0.19$
100-200 $4.97\pm 0.12$
200-300 $3.76\pm 0.18$ $1.57\pm 0.10$
300-500 $3.36\pm 0.21$ $1.10\pm 0.09$
500-800 $3.07\pm 0.31$ $1.18\pm 0.20$
$>$800 $2.75\pm 0.49$ $1.86\pm 0.33$
.3 One-dimensional distributions of ROIs and control regions
The one-dimensional significance distributions of our analyzed ROIs are presented in top panels of Fig. S3. The positive mean values of the significance in both regions indicate the detection of diffuse emission. As a comparison, the one-dimensional significance distributions in control regions, shifted along the Galactic latitudes by $20^{\circ}$, are shown in bottom panels. The results in control regions are consistent with the standard Gaussian distribution as expected for the background.
.4 Fluxes of diffuse emission measured by LHAASO
The diffuse emission fluxes of the inner and outer Galaxy regions measured by LHAASO, together with $1\sigma$ statistical errors and the estimated systematic uncertainties, are presented in Tables S2 and S3.
We compare the measured diffuse $\gamma$-ray fluxes with the total fluxes of CR electrons plus positrons in Fig. S4. Although the electron and positrons fluxes are higher than the diffuse $\gamma$-ray emission by a factor of several below several TeV, they are largely removed with our data-driven background estimate method as long as electrons and positrons do not concentrate in the Galactic plane as $\gamma$ rays.
Fig. S5 shows the fluxes of all emission in the ROIs without source masks.
Table S2: Diffuse emission fluxes in the inner Galaxy region measured by WCDA-KM2A, together with $1\sigma$ statistical uncertainties and systematic uncertainties.
$\langle E\rangle$ $\phi\pm\sigma_{\rm stat}\pm\sigma_{\rm sys}$
(TeV) (TeV-1 cm-2 s-1 sr-1)
1.20 $(2.07\pm 0.27\pm 0.49)\times 10^{-10}$
2.03 $(6.84\pm 0.47\pm 0.60)\times 10^{-11}$
3.54 $(1.57\pm 0.14\pm 0.11)\times 10^{-11}$
5.81 $(3.35\pm 0.37\pm 0.45)\times 10^{-12}$
10.55 $(7.52\pm 0.99\pm 0.99)\times 10^{-13}$
21.53 $(1.01\pm 0.20\pm 0.15)\times 10^{-13}$
31.62 $(4.21\pm 0.44\pm 0.49)\times 10^{-14}$
50.12 $(9.90\pm 0.97\pm 1.57)\times 10^{-15}$
79.43 $(2.35\pm 0.27\pm 0.34)\times 10^{-15}$
125.89 $(5.19\pm 0.67\pm 0.73)\times 10^{-16}$
199.53 $(1.41\pm 0.21\pm 0.12)\times 10^{-16}$
316.23 $(3.36\pm 0.72\pm 0.50)\times 10^{-17}$
501.19 $(6.48\pm 2.48\pm 1.84)\times 10^{-18}$
794.33 $(2.17\pm 0.92\pm 1.10)\times 10^{-18}$
Table S3: Diffuse emission fluxes in the outer Galaxy region measured by WCDA-KM2A, together with $1\sigma$ statistical uncertainties and systematic uncertainties.
$\langle E\rangle$ $\phi\pm\sigma_{\rm stat}\pm\sigma_{\rm sys}$
(TeV) (TeV-1 cm-2 s-1 sr-1)
3.23 $(8.72\pm 1.43\pm 2.20)\times 10^{-12}$
5.36 $(2.00\pm 0.36\pm 0.13)\times 10^{-12}$
9.91 $(3.58\pm 0.90\pm 0.44)\times 10^{-13}$
20.59 $(4.76\pm 1.65\pm 1.44)\times 10^{-14}$
31.62 $(2.03\pm 0.33\pm 0.22)\times 10^{-14}$
50.12 $(4.66\pm 0.76\pm 0.51)\times 10^{-15}$
79.43 $(7.81\pm 2.09\pm 1.18)\times 10^{-16}$
125.89 $(2.64\pm 0.52\pm 0.40)\times 10^{-16}$
199.53 $(8.60\pm 1.59\pm 2.99)\times 10^{-17}$
316.23 $(2.35\pm 0.54\pm 0.82)\times 10^{-17}$
501.19 $(4.67\pm 1.91\pm 1.64)\times 10^{-18}$
794.33 $(1.45\pm 0.69\pm 0.80)\times 10^{-18}$
Figure S4: Comparison of the diffuse $\gamma$-ray fluxes measured by LHAASO with the total electron plus positron fluxes by several experiments [25-27].
Figure S5: SEDs of all the $\gamma$-ray emission in the inner and outer regions without source masks. Error bars are the statistical errors.
.5 Comparison of longitude and latitude distributions with KM2A
Fig. S6 shows the comparison of longitude and latitude distributions of the WCDA measurements with the KM2A results. The trends of the data in the two different energy bands are similar with each other. Difference of the longitude distributions in some longitude bins (e.g., $65^{\circ}<l<85^{\circ}$) are likely due to different source masks in the two analyses.
Figure S6: Comparison of the latitude and longitude profiles of the WCDA (red points) and KM2A (blue points) measurements. The KM2A data for the $10-63$ TeV band are plotted [11].

### v3

@@ Abstract
The diffuse Galactic gamma-ray emission is a very important tool used to study the propagation and interaction of cosmic rays in the Milky Way. In this work, we report the measurements of the diffuse emission from the Galactic plane, covering Galactic longitudes from $15^{\circ}$ to $235^{\circ}$ and latitudes from $-5^{\circ}$ to $+5^{\circ}$, in an energy range of 1 TeV to 25 TeV, with the Water Cherenkov Detector Array (WCDA) of the Large High Altitude Air Shower Observatory (LHAASO). After masking the sky regions of known sources, the diffuse emission is detected with $24.6\sigma$ and $9.1\sigma$ significance in the inner Galactic plane($15^{\circ}<l<125^{\circ}$, $|b|<5^{\circ}$) and outer Galactic plane ($125^{\circ}<l<235^{\circ}$, $|b|<5^{\circ}$), respectively. The WCDA spectra in both regions can be well described by a power-law function, with spectral indices of $-2.67\pm 0.05_{\rm stat}$ in the inner region and $-2.83\pm 0.19_{\rm stat}$ in the outer region, respectively. Combined with the Square Kilometer Array (KM2A) measurements at higher energies, a clear softening of the spectrum is found in the inner region, with change of spectral indices by $\sim 0.5$ at a break energy around $30$ TeV. The fluxes of the diffuse emission are higher by a factor of $1.5-2.7$ than the model prediction assuming local CR spectra and the gas column density, which are consistent with those measured by the KM2A. Along Galactic longitude, the spatial distribution of the diffuse emission shows deviation from that of the gas column density. The spectral shape of the diffuse emission are possibly variation in different longitude region. The WCDA measurements bridge the gap between the low-energy measurements by space detectors and the ultra-high-energy observations by LHAASO-KM2A and other experiments. These results suggest that improved modeling of the wide-band diffuse emission is required.
@@
Introduction. — One of the most fundamental unresolved problems in astrophysics is the origin and propagation of cosmic rays (CRs). Besides the direct measurements of energy spectra and arrival directions of individual composition of CRs, the diffuse $\gamma$-ray emission plays a unique and complementary role in constraining the origin and propagation of CRs. The diffuse $\gamma$-ray emission is usually expected to be produced through interactions between CRs (nuclei and electrons/positrons) and the interstellar medium (ISM) or radiation field [1, 2, 3]. Modeling of the diffuse emission in the high-energy $\gamma$-ray bands (HE, $\lesssim$0.1 TeV) is consistent with the measurements at high and intermediate latitudes by Fermi Large Area Telescope [4], supporting the basic framework of CR propagation and interaction in the Milky Way.
@@
We also derive the spatial distributions of the WCDA diffuse emission along Galactic latitudes and longitudes, as shown in Fig. 4. The results are compared with the gas distributions obtained from the sum of atomic hydrogen from the HI4PI survey [30], molecular hydrogen from the CO survey [31] with a CO-to-H2 conversion factor of $X_{\rm CO}=1.6\times 10^{20}$ cm2 $({\rm K\,km\,s^{-1}})^{-1}$ [4], and an ionized hydrogen component from the model of [32]. An alternative way is to use the dust opacity to trace the gas column density [33]. Differences from these two methods vary for different regions of the Galaxy. On average, we find that the inferred gas column densities in the ROIs defined in this work are smaller than 10%. It is shown that the latitude distributions are roughly consistent with the gas distributions in both the inner and outer regions. The fitting of the gas template to the data gives $p$-values of $0.003-0.013$ of the two regions. However, the longitude distribution deviates clearly from the gas distribution. The gas template fitting shows a $p$-value of $6.0\times 10^{-21}$. Similar results have also been obtained for the KM2A measurements at higher energies [11] (see Fig. S6 of the Supplemental Material for the comparison between WCDA and KM2A).
@@
Systematic uncertainties. — One of the major sources of systematic uncertainties is the background estimate. We vary the time window from ±$\pm 5$ hours to ±$\pm 12$ hours, and also change the mask regions for background estimate to $|b|\leq 10^{\circ}$ for all declinations to obtain the variations of the results. The other effect related with the background is the spurious large-scale efficiency correction (see Sec. A of the Supplemental Material), for which we adopt different smoothing angles of $15^{\circ}$, $20^{\circ}$, $25^{\circ}$, and $30^{\circ}$. We find that the impacts of those changes on parameters $\phi_{0}$ are about $6\%~{}(16\%)$ and on $\alpha$ are about 0.05 (0.13), for the inner (outer) region. The variation of the atmosphere condition during the operation affects the detection efficiency, which is not properly modelled in the simulation. For point-like sources this effect is estimated to be about 8% for the flux ($\phi_{0}$) and 0.02 for the spectral index ($\alpha$) [19]. Similar impacts on the diffuse emission is expected. Combining these results, we obtain the total systematic uncertainties as 10% (18%) for $\phi_{0}$ and 0.05 (0.13) for $\alpha$, for the inner (outer) region, respectively. The discussion of systematic uncertainties of the KM2A results can be referred to Ref. [11]. The systematic uncertainty of the flux in each energy bin has been shown in Fig. 2 and also in Tables S2 and S3 of the Supplemental Material.
@@
Discussion. — We compare the data with the predictions of the hadronic interactions between CRs and the ISM. We adopt the parameterizations of the local proton and helium spectra as given in Ref. [11], which include the uncertainties of the current measurements, especially for those measured by indirect detection experiments above $\sim 100$ TeV. We assume that the CR intensity is uniform in the Galaxy111 Note that this assumption is over-simplified since variations of the CR intensities across the Galaxy were found to be present [34, 35]. Assuming a GALPROP model, which approximates the propagation halo as a cylinder, employs a numerical method to solve the transport equations, and includes experimental/observational inputs about the nuclear cross sections and gas distributions [2], we find that the expected diffuse fluxes are higher (lower) by 10% (40%) than the results obtained here for the inner (outer) region after the same source mask., and calculate the $\gamma$-ray emission using the local CR spectra and the line-of-sight integrated gas content. The $\gamma$-ray production cross section used is from the AAfrag package [36]. The expected diffuse $\gamma$-ray fluxes are shown by grey shaded bands in Fig. 2. We can see that in the inner region the measurements are higher by a factor of $\sim 2.7$ than the prediction, and in the outer region the measurements are higher by $\sim 1.5$ times. These results are consistent with those found in the LHAASO-KM2A analysis [11]. The HAWC measurements in a different sky region of the inner Galaxy also showed a factor of 2 higher fluxes compared with a DRAGON model prediction [10]. In the outer region, the predicted spectrum is consistent with the data within uncertainties. We test the consistency of the spectral shape between the prediction and the measurement in the inner region through a fit with free normalization of the model curve, and get $\chi^{2}=86.14$ and 19.16 for 13 degrees of freedom, for the “high” and “low” predictions (corresponding to the upper and lower edges of the band shown in Fig. 2). The $p$-values of the fittings are $7.61\times 10^{-13}$ and 0.12, respectively. Given the current large uncertainties of the spectral measurements of individual CR species, we cannot draw a firm conclusion on whether the diffuse $\gamma$-ray data is consistent with the knee of the light CR components or not based on the spectral shape only.
@@
If we compare the data (including those measured by Fermi-LAT between 1 and 500 GeV [12], updated with the new mask of this work) with the GALPROP model prediction222 The propagation and source injection parameters were adjusted [37] to fit the up-to-date measurements of CR primary and secondary nuclei [38, 39, 40]. One should be aware that, a homogeneous, isotropic diffusion is assumed when calculating the CR propagation, which may be too simple [41]. in a wider energy range, from GeV to PeV, we find that the excess mainly exists from several GeV to tens of TeV (see Fig. 5). Such an excess may come from a population of unresolved, low surface brightness sources such as pulsar wind nebulae or pulsar halos [1, 42, 43, 13, 16]. The spatial variations of the spectra of the diffuse emission may further supports the hypothesis that unresolved sources may exist, although other possibilities such as spatially-dependent spectral slopes of CRs can not be excluded. Following Ref. [12], we add an empirical component with spectrum $\propto E^{-2.45}\exp\left(-E/20\,{\rm TeV}\right)$ to the CR propagation model prediction, and find a good agreement with the wide-band measurements, as shown in Fig. 5. Other sources like young massive star clusters [44], the confinement and interaction of CRs around acceleration sources [45, 46, 47, 48], as well as the change of the canonical propagation scenario [49, 50, 51, 52] were also discussed in literature. The multi-messenger analysis of the diffuse $\gamma$ rays and neutrinos [29] would be helpful in understanding the nature of the diffuse emission.
@@
Summary. — In this work we measure the diffuse $\gamma$-ray emission in an energy range of about $1-25$ TeV from the Galactic plane using the LHAASO-WCDA data. The detection significance of the diffuse emission, after masking sky regions around resolved sources, is $24.6\sigma$ for the inner Galactic plane ($15^{\circ}<l<125^{\circ}$, $|b|<5^{\circ}$) and $9.1\sigma$ for the outer Galactic plane ($125^{\circ}<l<235^{\circ}$, $|b|<5^{\circ}$) regions, respectively. The diffuse emission from the outer Galactic plane is for the first time detected in the multi-TeV energy range. The WCDA measurements fill the gap between direct detection by space telescopes and ultra-high-energy measurements by groundbased experiments, and offer a full energy coverage from GeV to PeV for the Galactic diffuse $\gamma$-ray emission. The SEDs of the diffuse emission in the two regions have been measured with high precision. The WCDA spectra are consistent with power-law models, with indices of $-2.67\pm 0.05_{\rm stat}$ and $-2.83\pm 0.19_{\rm stat}$. The joint WCDA-KM2A spectrum shows a break around 30 TeV in the inner region, with indices changing by about $0.5$. The break in the outer region is not significant. We also show that there are slight spectral variations across the Galactic plane in the inner region. The measured fluxes are higher by a factor of $1.5-2.7$ than the simple prediction of hadronic interactions between CRs (with locally measured fluxes) and the ISM, indicating that there are unmodelled source components or spatial variations of CR distribution.
@@
Acknowledgements. — We would like to thank all staff members who work at the LHAASO site above 4400 meter above the sea level year-round to maintain the detector and keep the water recycling system, electricity power supply and other components of the experiment operating smoothly. We are grateful to Chengdu Management Committee of Tianfu New Area for the constant financial support for research with LHAASO data. We appreciate the computing and data service support provided by the National High Energy Physics Data Center for the data analysis in this paper. This work is supported by the following grants: the Department of Science and Technology of Sichuan Province, China (No. 2024ZYD0111, 24NSFSC2319), the National Natural Science Foundation of China (Nos. 12220101003, 12173039, 12393851, 12393852, 12393853, 12393854, 12205314, 12105301, 12305120, 12261160362, 12105294, U1931201, 12375107, 12342502), the Project for Young Scientists in Basic Research of Chinese Academy of Sciences (No. YSBR-061), the National Science and Technology Development Agency (NSTDA) of Thailand, and the National Research Council of Thailand (NRCT) under the High-Potential Research Team Grant Program (N42A650868).